What is the column percentage for "Package design C" and "Age 25-40"?

A market research firm has conducted a study to determine consumer preference for a new package design for a particular product. The consumers, ages were also noted 

            Package Design
AgeABCTotal
Under 2518182965
25-401812535
Total363034100

What is the column percentage for "Package design C" and "Age 25-40"? 

a. 34.29% 

b. 50% 

c. 14.71% 

d. 51.43%
Answer #1

Since it's a case of column percentage case so simply take corresponding value which results from " Package Design C" and "Age 25-40 " :

the value is : 5

and divide it by total number of "Package Design C" :

the value is : 34

So the answer will be = (5) / (34) = 0.1471

In percentage = 0.1472 * 100 = 14.71 %

The answer is option "C"